what is the 7nth root of -2187a raised to 8
Answer by Alan3354(69443)

3^7 = 2187
The 7th root is either 3 or -3, depending on whether the minus sign is inside or outside the 7th root.
Either way, 3^8 = 6561
and, (-3)^8 = 6561
-----------------------
I don't know what the "a" is in 2187a, I ignored it.